Use the information in the table below:

Year Nominal GDP Real GDP

2012 280 290

2013 315 260

2014 305 310

  1. The increase in nominal GDP between 2012 and 2014 is:

1.1. -3.5 %

1.2. 19.2 %

1.3. 8.93 %

1.4. 6.90 %


2.The increase in real GDP between 2012 and 2014 is:

2.1. 5 %

2.2. 6.90 %

2.3. 7.14 %

2.4. 20 %


1
Expert's answer
2021-09-28T13:18:23-0400

Solution:

1.). The increase in nominal GDP between 2012 and 2014 is 1.4.) 8.93"\%"

Nominal GDP in 2012 = 280

Nominal GDP in 2014 = 305

Percentage increase = "\frac{New\; value - Old \;value}{Old \;value}\times 100\%"


= "\frac{305 - 280}{280}\times 100\% = 8.93\%"


The increase in nominal GDP between 2012 and 2014 is 1.4.) 8.93"\%"

 

2.). The increase in real GDP between 2012 and 2014 is 1.4.) 6.90"\%"

Real GDP in 2012 = 290

Real GDP in 2014 = 310

Percentage increase = "\frac{New\; value - Old \;value}{Old \;value}\times 100\%"


= "\frac{310 - 290}{290}\times 100\% = 6.90\%"


The increase in real GDP between 2012 and 2014 is 2.2.) 6.90"\%"